OpenStudy (anonymous):

Write an inequality to represent the problem, and then solve the inequality by writing the pairs which solve it . Find all sets of two consecutive positive odd integers whose sum is less than or equal to 18.

OpenStudy (anonymous):

\[n+n+1\leq 18\] \[2n+1\leq 18\] \[2n\leq 17\] \[n\leq 8\tfrac{1}{2}\]

OpenStudy (anonymous):

so maybe 1,2 2,3 3,4 4,5 5,6 .....

OpenStudy (anonymous):

thank you so much